# Renewal theory for Brownian motion with stochastically gated targets

### v1

@@ 3.1 Brownian particle with stochastic resetting on the half-line
Let us first return to the example of a Brownian particle on a half-line with a stochastically gated boundary at $x=0$. However, we now assume that when the particle is diffusing in the bulk domain $(0,\infty)$ it instantaneously resets to a fixed reset position $x_{r}$ at a random sequence of times generated from a Poisson process with constant rate $r$ [49, 50]. (We suppress the explicit dependence on $x_{r}$. At the end of the calculation we will set $x_{r}=x_{0}$.) The first renewal equation (2.17) still holds except that the forward Kolmogorov equation for $p$ becomes
$\frac{\partial p(x,t|x_{0})}{\partial t}$ $=D\frac{\partial^{2}p(x,t|x_{0})}{\partial x^{2}}-rp(x,t|x_{0})+rQ(x_{0},t)\delta(x-x_{r}),$ (3.1a)
$p(0,t|x_{0})$ $=0,\quad p(x,0|x_{0})=\delta(x-x_{0}),$ (3.1b)
@@
$Q(x_{0},t)=\int_{0}^{\infty}p(x,t|x_{0})dx.$ (3.2)
Equation (3.1) can either be solved directly using Laplace transforms or by using renewal theory [49, 50]. Here we use the former approach. Laplace transforming equations (3.1a,b) gives
$D\frac{\partial^{2}\widetilde{p}(x,s|x_{0})}{\partial x^{2}}-(r+s)\widetilde{p}(x,s|x_{0})=-\delta(x-x_{0})-r\widetilde{Q}(x_{0},s)\delta(x-x_{r}),\quad x>0,$ (3.3)
@@
$\widetilde{p}(x,s|x_{0})$ $=\widetilde{G}(x,r+s|x_{0})+r\widetilde{Q}(x_{0},s)\widetilde{G}(x,r+s|x_{r}),$ (3.4)
with $\widetilde{G}(x,s|x_{0})$ given by equation (2.10). Finally, Laplace transforming equation (3.2) and using (3.4) shows that
$\widetilde{Q}(x_{0},s)$ $=\int_{0}^{\infty}\widetilde{p}(x,s|x_{0})dx=\widetilde{Q}_{0}(x_{0},r+s)+r\widetilde{Q}(x_{0},s)\widetilde{Q}_{0}(x_{r},r+s),$ (3.5)
where $\widetilde{Q}_{0}(x_{0},s)$ is the survival probability without resetting, see equation (2.33). Rearranging equation (3.5) then gives the well-known results [49, 50]
$\widetilde{Q}(x_{0},s)$ $=\frac{\widetilde{Q}_{0}(x_{0},r+s)}{1-r\widetilde{Q}_{0}(x_{r},r+s)},$ (3.6a)
$\widetilde{f}(x_{0},s)$ $\equiv\frac{1-(r+s)\widetilde{Q}_{0}(x_{0},r+s)}{1-r\widetilde{Q}_{0}(x_{r},r+s)}+\frac{r\bigg[\widetilde{Q}_{0}(x_{0},r+s)-\widetilde{Q}_{0}(x_{r},r+s)\bigg]}{1-r\widetilde{Q}_{0}(x_{r},r+s)}.$ (3.6b)
In order to take the limit $\epsilon\rightarrow 0$ in equation (2.30) we need to determine the leading order contributions to $\widetilde{Q}(\epsilon,s)$ and $\widetilde{f}(\epsilon,s)$. Hence, setting $x_{0}=\epsilon$ in equations (3.6a,b), we have
$\widetilde{Q}(\epsilon,s)$ $=\frac{\epsilon/\sqrt{[r+s]D}}{1-r\widetilde{Q}_{0}(x_{r},r+s)}+O(\epsilon^{2}),\$(3.7a)
@@
$\widetilde{f}(\epsilon,s)$ $\equiv\frac{1-\epsilon\sqrt{[r+s]/D}}{1-r\widetilde{Q}_{0}(x_{r},r+s)}+\frac{r\bigg[\epsilon/\sqrt{[r+s]D}-\widetilde{Q}_{0}(x_{r},r+s)\bigg]}{1-r\widetilde{Q}_{0}(x_{r},r+s)}+O(\epsilon^{2})$
$=1-\frac{\epsilon}{1-r\widetilde{Q}_{0}(x_{r},r+s)}\frac{s}{\sqrt{[r+s]D}}+O(\epsilon^{2}).$ (3.7b)
Combining equations (2.24) and (2.34) then yields
$\lim_{\epsilon\rightarrow 0}\ \frac{\widetilde{Q}(\epsilon,s)}{1-\widetilde{f}_{-1}(\epsilon,s)}$ (3.8)
@@
Now taking the limit $s\rightarrow 0$ in equation (2.30) and setting $x_{r}=x_{0}$ yields the MFPT
${\mathcal{T}}_{k_{0}}(x_{0})$ $=\frac{\widetilde{Q}_{0}(x_{0},r)}{1-r\widetilde{Q}_{0}(x_{0},r)}+\widetilde{f}_{k_{0}}(x_{0},0)\frac{{\mathrm{e}}^{\sqrt{r/D}x_{0}}}{\sqrt{r}}\frac{\sqrt{r+s+2\alpha}}{\alpha}\bigg[1-r\widetilde{Q}_{0}(x_{0},r+2\alpha)\bigg],$ (3.9)
@@
$\widetilde{f}_{k_{0}}(x_{0},0)=\frac{1}{2}\left(1-k_{0}\frac{1-(r+2\alpha)\widetilde{Q}_{0}(x_{0},r+2\alpha)}{1-r\widetilde{Q}_{0}(x_{0},r+2\alpha)}\right).$ (3.10)
@@
${\mathcal{T}}_{k_{0}}(x_{0})$ $=\frac{\widetilde{Q}_{0}(x_{0},r)}{1-r\widetilde{Q}_{0}(x_{0},r)}$
$\quad+\frac{{\mathrm{e}}^{\sqrt{r/D}x_{0}}}{\sqrt{r}}\frac{\sqrt{r+2\alpha}}{2\alpha}\bigg[1-k_{0}+[k_{0}(r+2\alpha)-r]\widetilde{Q}_{0}(x_{0},r+2\alpha)\bigg].$ (3.11)
It can be checked that equation (3.1) recovers the MFPT for $k_{0}=\pm 1$ derived in Ref. [51] by directly solving the forward Kolmogorov equation for $\rho_{k|k_{0}}$ in the presence of resetting. Also note that in the slow switching limit, we have
$\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{-1}(x_{0})=\infty,\quad\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{1}(x_{0})=\frac{2\widetilde{Q}_{0}(x_{0},r)}{1-r\widetilde{Q}_{0}(x_{0},r)}=2\frac{{\mathrm{e}}^{\sqrt{r/D}x_{0}}-1}{r}.$ (3.12)
@@
One useful feature of the renewal formulation is that we can straightforwardly generalise the model by, for example, keeping $\epsilon>0$ fixed rather than taking the limit $\epsilon\rightarrow 0$ in equation (2.30). (In Ref. [46] we present an argument for a particle resetting to its initial position ($\epsilon=x_{0}$) when leaving a boundary without being absorbed, based on a foraging animal returning to its homebase after failing to capture resources.) Therefore, consider
$\widetilde{\mathfrak{Q}}_{k_{0}}(x_{0},s)=\widetilde{Q}(x_{0},s)+\widetilde{f}_{k_{0}}(x_{0},s)\frac{\widetilde{Q}(\epsilon,s)}{1-\widetilde{f}_{-1}(\epsilon,s)},\quad\epsilon>0.$ (3.13)
with $\widetilde{f}_{k_{0}}(x_{0},s)$ defined by equations (2.24) and (3.6). In Fig. 5 we show example plots of the corresponding MFPT ${\mathcal{T}}_{k_{0}}(x_{0})=\lim_{s\rightarrow 0}\widetilde{\mathfrak{Q}}_{k_{0}}(x_{0},s)$ as a function of the resetting rate $r$ for different values of $\epsilon$ and $\alpha$. It can be seen that the solutions converge in the limit $\epsilon\rightarrow 0$. A number of other features can be observed. First, each curve has the characteristic unimodal shape indicating the existence of an optimal resetting rate for minimising the MFPT. Second, ${\mathcal{T}}_{1}(x_{0})<{\mathcal{T}}_{-1}(x_{0})$ for the same set of parameter values, that is, the MFPT is reduced if the gate is initially open. Third, reducing the switching rate $\alpha$ increases the gap between ${\mathcal{T}}_{1}(x_{0})$ and ${\mathcal{T}}_{-1}(x_{0})$ with $\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{-1}(x_{0})=\infty$.
@@ 3.2.1 Correlated gates
Let us first consider the totally correlated case with $\Pi_{kk_{0}}(t)$ specifying the probability that the particle is in the state $k$ at time $t$, given that it was initially in state $k_{0}$. Using the casual structure of the renewal formulation, it is straightforward to write down the analogue of the first renewal equation (2.17):
$\rho^{(\epsilon)}_{k|k_{0}}(x,t|x_{0})$ $=\Pi_{kk_{0}}(t)p(x,t|x_{0})-\int_{0}^{t}\rho^{(\epsilon)}_{k|-1}(x,t-\tau|\epsilon)\Pi_{-1,k_{0}}(\tau)J(0,\tau|x_{0})d\tau$
$\quad+\int_{0}^{t}\rho^{(\epsilon)}_{k|-1}(x,t-\tau|L-\epsilon)\Pi_{-1,k_{0}}(\tau)J(L,\tau|x_{0})d\tau.$ (3.14)
@@
$\frac{\partial p(x,t|x_{0})}{\partial t}$ $=D\frac{\partial^{2}p(x,t|x_{0})}{\partial x^{2}},\quad 0<x<L,$ (3.15a)
$p(0,t|x_{0})$ $=0=p(L,t|x_{0}),\quad p(x,0|x_{0})=\delta(x-x_{0}),$ (3.15b)
@@
$\pi_{L}(x_{0})=\int_{0}^{\infty}J(L,t|x_{0})dt,\quad f_{L}(x_{0},t)=\frac{J(L,t|x_{0})}{\pi_{L}(0)}.$ (3.17)
@@
$\pi_{0}(x_{0})=-\int_{0}^{\infty}J(0,t|x_{0})dt,\quad f_{0}(x_{0},t)=-\frac{J(0,t|x_{0})}{\pi_{0}}.$ (3.18)
Laplace transforming equation (3.2.1) we have
$\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(x,s|x_{0})$ $=\widetilde{p}_{k|k_{0}}(x,s|x_{0})-\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\widetilde{J}_{k_{0}}(0,s|x_{0})$ (3.19)
@@
where $\widetilde{p}_{k|k_{0}}(x,s|x_{0})$ is given by equation (2.2) and
$\widetilde{J}_{k_{0}}(x,s|x_{0})$ $=\frac{1}{2}\bigg(\widetilde{J}(x,s|x_{0})-k_{0}\widetilde{J}(x,s+2\alpha|x_{0})\bigg).$ (3.20)
In order to eliminate the terms $\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)$ and $\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)$ on the right-hand side of equation (3.2.1) we set $k_{0}=-1$ and either $x_{0}=\epsilon$ or $x_{0}=L-\epsilon$. This leads to the pair of simultaneous equations
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)$ $=\widetilde{p}_{k|-1}(x,s|\epsilon)-\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\widetilde{J}_{-1}(0,s|\epsilon)$
$\quad+\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|L-\epsilon)\widetilde{J}_{-1}(L,s|\epsilon),$ (3.21a)
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|L-\epsilon)$ $=\widetilde{p}_{k|-1}(x,s|L-\epsilon)-\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\widetilde{J}_{-1}(0,s|L-\epsilon)$
$\quad+\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|L-\epsilon)\widetilde{J}_{-1}(L,s|L-\epsilon).$ (3.21b)
@@
$\left(\begin{array}[]{c}\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\\ \widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|L-\epsilon)\end{array}\right)=\left[{\bf I}-{\bf M}(\epsilon,s)\right]^{-1}\left(\begin{array}[]{c}\widetilde{p}_{k|-1}(x,s|\epsilon)\\ \widetilde{p}_{k|-1}(x,s|L-\epsilon)\end{array}\right),$ (3.26)
@@
${\bf M}(\epsilon,s):=\left(\begin{array}[]{cc}-\widetilde{J}_{-1}(0,s|\epsilon)&\widetilde{J}_{-1}(L,s|\epsilon)\\ -\widetilde{J}_{-1}(0,s|L-\epsilon)&\widetilde{J}_{-1}(L,s|L-\epsilon)\end{array}\right).$ (3.29)
Substituting the solution (3.26) into equation (3.19) and taking the limit $\epsilon\rightarrow 0$ gives
$\widetilde{\rho}_{k|k_{0}}(x,s|x_{0})$ $=\widetilde{p}_{k|k_{0}}(x,s|x_{0})$ (3.30)
@@
$\widetilde{{\mathfrak{Q}}}_{k_{0}}(x_{0},s)$ $=\widetilde{Q}(x_{0},s)$ (3.31)
$\quad-\lim_{\epsilon\rightarrow 0}\sum_{a,b=0,1}(-1)^{a}\widetilde{J}_{k_{0}}(aL,s|x_{0})\left[{\bf I}-{\bf M}(\epsilon,s)\right]_{ab}^{-1}\widetilde{Q}(|bL-\epsilon|,s).$
@@
${\mathfrak{Q}}_{k_{0}}(x_{0},s)$ $:=\int_{0}^{L}\sum_{k=\pm 1}\widetilde{\rho}_{k|k_{0}}(x,s|x_{0})dx,$ (3.32a)
$\widetilde{Q}(x_{0},s)$ $:=\int_{0}^{L}\widetilde{p}(x,s|x_{0})dx=\frac{1-\widetilde{J}(L,s|x_{0})+\widetilde{J}(0,s|x_{0})}{s}.$ (3.32b)
Equation (3.31) is the 2-gate analogue of equation (2.30).
@@
$G(x,s|x_{0})$ $=\frac{\sinh\sqrt{s/D}x\,\sinh\sqrt{s/D}(L-x_{0})}{\sqrt{sD}\sinh\sqrt{s/D}L}\Theta(x_{0}-x)$ (3.33)
$\quad+\frac{\sinh\sqrt{s/D}x_{0}\,\sinh\sqrt{s/D}(L-x)}{\sqrt{sD}\sinh\sqrt{s/D}L}\Theta(x-x_{0}),$ (3.34)
@@
$\widetilde{J}(0,s|x_{0})$ $=-\frac{\sinh\sqrt{s/D}(L-x_{0})}{\sinh\sqrt{s/D}L},\quad\widetilde{J}(L,s|x_{0})=\frac{\sinh\sqrt{s/D}x_{0}}{\sinh\sqrt{s/D}L}.$ (3.35)
Substituting into the matrix (3.39) and expanding to $O(\epsilon)$ gives
${\bf M}(\epsilon,s)={\bf I}-\epsilon{\bm{\Lambda}}(s)+O(\epsilon^{2}),$ (3.36)
@@
${\bf m}(s)=\frac{\sqrt{{s}/{D}}}{2\sinh\sqrt{s/D}L}\left(\begin{array}[]{cc}\mbox{cosh}\sqrt{s/D}L&-1\\ -1&\mbox{cosh}\sqrt{s/D}L\end{array}\right).$ (3.39)
@@
$\widetilde{Q}(L-\epsilon,s)=\widetilde{Q}(\epsilon,s)=\epsilon\widetilde{q}(s)+O(\epsilon^{2}),\quad\widetilde{q}(s)\equiv\frac{\cosh\sqrt{s/D}L-1}{\sqrt{sD}\sinh\sqrt{s/D}L},$ (3.40)

### v2

@@ 3.1 Brownian particle with stochastic resetting on the half-line
Let us first return to the example of a Brownian particle on a half-line with a stochastically gated boundary at $x=0$. However, we now assume that when the particle is diffusing in the bulk domain $(0,\infty)$ it instantaneously resets to a fixed reset position $x_{r}$ at a random sequence of times generated from a Poisson process with constant rate $r$ [49, 50]. (We suppress the explicit dependence on $x_{r}$. At the end of the calculation we will set $x_{r}=x_{0}$.) The first renewal equation (2.4gr) still holds except that the forward Kolmogorov equation for $p$ becomes
$\frac{\partial p(x,t|x_{0})}{\partial t}=D\frac{\partial^{2}p(x,t|x_{0})}{\partial x^{2}}-rp(x,t|x_{0})+rQ(x_{0},t)\delta(x-x_{r}),$ (2.4gsajaa)
$p(0,t|x_{0})=0,\quad p(x,0|x_{0})=\delta(x-x_{0}),$ (2.4gsajab)
@@
$Q(x_{0},t)=\int_{0}^{\infty}p(x,t|x_{0})dx.$ (2.4gsajab)
Equations (2.4gsajaa) and (2.4gsajab) can either be solved directly using Laplace transforms or by using renewal theory [49, 50]. Here we use the former approach. Laplace transforming equations (2.4gsajaa) and (2.4gsajab) gives
$D\frac{\partial^{2}\widetilde{p}(x,s|x_{0})}{\partial x^{2}}-(r+s)\widetilde{p}(x,s|x_{0})=-\delta(x-x_{0})-r\widetilde{Q}(x_{0},s)\delta(x-x_{r}),\quad x>0,$ (2.4gsajac)
@@
$\widetilde{p}(x,s|x_{0})$ $=\widetilde{G}(x,r+s|x_{0})+r\widetilde{Q}(x_{0},s)\widetilde{G}(x,r+s|x_{r}),$ (2.4gsajad)
with $\widetilde{G}(x,s|x_{0})$ given by equation (2.4gk). Finally, Laplace transforming equation (2.4gsajab) and using (2.4gsajad) shows that
$\widetilde{Q}(x_{0},s)$ $=\int_{0}^{\infty}\widetilde{p}(x,s|x_{0})dx=\widetilde{Q}_{0}(x_{0},r+s)+r\widetilde{Q}(x_{0},s)\widetilde{Q}_{0}(x_{r},r+s),$ (2.4gsajae)
where $\widetilde{Q}_{0}(x_{0},s)$ is the survival probability without resetting, see equation (2.4gsah). Rearranging equation (2.4gsajae) then gives the well-known results [49, 50]
$\widetilde{Q}(x_{0},s)$ $=\frac{\widetilde{Q}_{0}(x_{0},r+s)}{1-r\widetilde{Q}_{0}(x_{r},r+s)},$ (2.4gsajafa)
$\widetilde{f}(x_{0},s)$ $\equiv\frac{1-(r+s)\widetilde{Q}_{0}(x_{0},r+s)}{1-r\widetilde{Q}_{0}(x_{r},r+s)}+\frac{r\bigg[\widetilde{Q}_{0}(x_{0},r+s)-\widetilde{Q}_{0}(x_{r},r+s)\bigg]}{1-r\widetilde{Q}_{0}(x_{r},r+s)}.$ (2.4gsajafb)
In order to take the limit $\epsilon\rightarrow 0$ in equation (2.4gsae) we need to determine the leading order contributions to $\widetilde{Q}(\epsilon,s)$ and $\widetilde{f}(\epsilon,s)$. Hence, setting $x_{0}=\epsilon$ in equations (2.4gsajafa) and (2.4gsajafb), we have
$\widetilde{Q}(\epsilon,s)$ $=\frac{\epsilon/\sqrt{[r+s]D}}{1-r\widetilde{Q}_{0}(x_{r},r+s)}+O(\epsilon^{2}),\$(2.4gsajafga)
@@
$\widetilde{f}(\epsilon,s)$ $\equiv\frac{1-\epsilon\sqrt{[r+s]/D}}{1-r\widetilde{Q}_{0}(x_{r},r+s)}+\frac{r\bigg[\epsilon/\sqrt{[r+s]D}-\widetilde{Q}_{0}(x_{r},r+s)\bigg]}{1-r\widetilde{Q}_{0}(x_{r},r+s)}+O(\epsilon^{2})$ (2.4gsajafgb)
$=1-\frac{\epsilon}{1-r\widetilde{Q}_{0}(x_{r},r+s)}\frac{s}{\sqrt{[r+s]D}}+O(\epsilon^{2}).$
Combining with equation (2.4gsy) then yields
$\lim_{\epsilon\rightarrow 0}\ \frac{\widetilde{Q}(\epsilon,s)}{1-\widetilde{f}_{-1}(\epsilon,s)}$ (2.4gsajafgh)
@@
Now taking the limit $s\rightarrow 0$ in equation (2.4gsae) and setting $x_{r}=x_{0}$ yields the MFPT
${\mathcal{T}}_{k_{0}}(x_{0})$ $=\frac{\widetilde{Q}_{0}(x_{0},r)}{1-r\widetilde{Q}_{0}(x_{0},r)}+\widetilde{f}_{k_{0}}(x_{0},0)\frac{{\mathrm{e}}^{\sqrt{r/D}x_{0}}}{\sqrt{r}}\frac{\sqrt{r+s+2\alpha}}{\alpha}\bigg[1-r\widetilde{Q}_{0}(x_{0},r+2\alpha)\bigg],$
@@
$\widetilde{f}_{k_{0}}(x_{0},0)=\frac{1}{2}\left(1-k_{0}\frac{1-(r+2\alpha)\widetilde{Q}_{0}(x_{0},r+2\alpha)}{1-r\widetilde{Q}_{0}(x_{0},r+2\alpha)}\right).$ (2.4gsajafgj)
@@
${\mathcal{T}}_{k_{0}}(x_{0})$ $=\frac{\widetilde{Q}_{0}(x_{0},r)}{1-r\widetilde{Q}_{0}(x_{0},r)}$ (2.4gsajafgk)
$\quad+\frac{{\mathrm{e}}^{\sqrt{r/D}x_{0}}}{\sqrt{r}}\frac{\sqrt{r+2\alpha}}{2\alpha}\bigg[1-k_{0}+[k_{0}(r+2\alpha)-r]\widetilde{Q}_{0}(x_{0},r+2\alpha)\bigg].$
It can be checked that equation (2.4gsajafgk) recovers the MFPT for $k_{0}=\pm 1$ derived in Ref. [51] by directly solving the forward Kolmogorov equation for $\rho_{k|k_{0}}$ in the presence of resetting. Also note that in the slow switching limit, we have
$\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{-1}(x_{0})=\infty,\quad\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{1}(x_{0})=\frac{2\widetilde{Q}_{0}(x_{0},r)}{1-r\widetilde{Q}_{0}(x_{0},r)}=2\frac{{\mathrm{e}}^{\sqrt{r/D}x_{0}}-1}{r}.$ (2.4gsajafgl)
@@
One useful feature of the renewal formulation is that we can straightforwardly generalise the model by, for example, keeping $\epsilon>0$ fixed rather than taking the limit $\epsilon\rightarrow 0$ in equation (2.4gsae). (In Ref. [46] we present an argument for a particle resetting to its initial position ($\epsilon=x_{0}$) when leaving a boundary without being absorbed, based on a foraging animal returning to its homebase after failing to capture resources.) Therefore, consider
$\widetilde{\mathfrak{Q}}_{k_{0}}(x_{0},s)=\widetilde{Q}(x_{0},s)+\widetilde{f}_{k_{0}}(x_{0},s)\frac{\widetilde{Q}(\epsilon,s)}{1-\widetilde{f}_{-1}(\epsilon,s)},\quad\epsilon>0,$ (2.4gsajafgm)
with $\widetilde{f}_{k_{0}}(x_{0},s)$ defined by equations (2.4gsy) and (2.4gsajafb). In Fig. 5 we show example plots of the corresponding MFPT ${\mathcal{T}}_{k_{0}}(x_{0})=\lim_{s\rightarrow 0}\widetilde{\mathfrak{Q}}_{k_{0}}(x_{0},s)$ as a function of the resetting rate $r$ for different values of $\epsilon$ and $\alpha$. It can be seen that the solutions converge in the limit $\epsilon\rightarrow 0$. A number of other features can be observed. First, each curve has the characteristic unimodal shape indicating the existence of an optimal resetting rate for minimising the MFPT. Second, ${\mathcal{T}}_{1}(x_{0})<{\mathcal{T}}_{-1}(x_{0})$ for the same set of parameter values, that is, the MFPT is reduced if the gate is initially open. Third, reducing the switching rate $\alpha$ increases the gap between ${\mathcal{T}}_{1}(x_{0})$ and ${\mathcal{T}}_{-1}(x_{0})$ with $\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{-1}(x_{0})=\infty$.
@@ 3.2.1 Correlated gates
Let us first consider the totally correlated case with $\Pi_{kk_{0}}(t)$ specifying the probability that the particle is in the state $k$ at time $t$, given that it was initially in state $k_{0}$. Using the casual structure of the renewal formulation, it is straightforward to write down the analogue of the first renewal equation (2.4gr):
$\rho^{(\epsilon)}_{k|k_{0}}(x,t|x_{0})$ $=\Pi_{kk_{0}}(t)p(x,t|x_{0})-\int_{0}^{t}\rho^{(\epsilon)}_{k|-1}(x,t-\tau|\epsilon)\Pi_{-1,k_{0}}(\tau)J(0,\tau|x_{0})d\tau$ (2.4gsajafgn)
$\quad+\int_{0}^{t}\rho^{(\epsilon)}_{k|-1}(x,t-\tau|L-\epsilon)\Pi_{-1,k_{0}}(\tau)J(L,\tau|x_{0})d\tau.$
@@
$\frac{\partial p(x,t|x_{0})}{\partial t}=D\frac{\partial^{2}p(x,t|x_{0})}{\partial x^{2}},\quad 0<x<L,$ (2.4gsajafgoa)
$p(0,t|x_{0})=0=p(L,t|x_{0}),\quad p(x,0|x_{0})=\delta(x-x_{0}),$ (2.4gsajafgob)
@@
$\pi_{L}(x_{0})=\int_{0}^{\infty}J(L,t|x_{0})dt,\quad f_{L}(x_{0},t)=\frac{J(L,t|x_{0})}{\pi_{L}(0)}.$ (2.4gsajafgoq)
@@
$\pi_{0}(x_{0})=-\int_{0}^{\infty}J(0,t|x_{0})dt,\quad f_{0}(x_{0},t)=-\frac{J(0,t|x_{0})}{\pi_{0}}.$ (2.4gsajafgor)
Laplace transforming equation (2.4gsajafgn) we have
$\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(x,s|x_{0})$ $=\widetilde{p}_{k|k_{0}}(x,s|x_{0})-\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\widetilde{J}_{k_{0}}(0,s|x_{0})$ (2.4gsajafgos)
@@
where $\widetilde{p}_{k|k_{0}}(x,s|x_{0})$ is given by equation (2.4gsx) and
$\widetilde{J}_{k_{0}}(x,s|x_{0})$ $=\frac{1}{2}\bigg(\widetilde{J}(x,s|x_{0})-k_{0}\widetilde{J}(x,s+2\alpha|x_{0})\bigg).$ (2.4gsajafgot)
In order to eliminate the terms $\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)$ and $\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)$ on the right-hand side of equation (2.4gsajafgos) we set $k_{0}=-1$ and either $x_{0}=\epsilon$ or $x_{0}=L-\epsilon$. This leads to the pair of simultaneous equations
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)$ $=\widetilde{p}_{k|-1}(x,s|\epsilon)-\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\widetilde{J}_{-1}(0,s|\epsilon)$ (2.4gsajafgoua)
$\quad+\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|L-\epsilon)\widetilde{J}_{-1}(L,s|\epsilon),$
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|L-\epsilon)$ $=\widetilde{p}_{k|-1}(x,s|L-\epsilon)-\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\widetilde{J}_{-1}(0,s|L-\epsilon)$ (2.4gsajafgoub)
$\quad+\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|L-\epsilon)\widetilde{J}_{-1}(L,s|L-\epsilon).$
@@
$\left(\begin{array}[]{c}\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\\ \widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|L-\epsilon)\end{array}\right)=\left[{\bf I}-{\bf M}(\epsilon,s)\right]^{-1}\left(\begin{array}[]{c}\widetilde{p}_{k|-1}(x,s|\epsilon)\\ \widetilde{p}_{k|-1}(x,s|L-\epsilon)\end{array}\right),$ (2.4gsajafgouz)
@@
${\bf M}(\epsilon,s):=\left(\begin{array}[]{cc}-\widetilde{J}_{-1}(0,s|\epsilon)&\widetilde{J}_{-1}(L,s|\epsilon)\\ -\widetilde{J}_{-1}(0,s|L-\epsilon)&\widetilde{J}_{-1}(L,s|L-\epsilon)\end{array}\right).$ (2.4gsajafgouac)
Substituting the solution (2.4gsajafgouz) into equation (2.4gsajafgos) and taking the limit $\epsilon\rightarrow 0$ gives
$\widetilde{\rho}_{k|k_{0}}(x,s|x_{0})$ $=\widetilde{p}_{k|k_{0}}(x,s|x_{0})$
@@
$\widetilde{{\mathfrak{Q}}}_{k_{0}}(x_{0},s)$ $=\widetilde{Q}(x_{0},s)-\lim_{\epsilon\rightarrow 0}\sum_{a,b=0,1}(-1)^{a}\widetilde{J}_{k_{0}}(aL,s|x_{0})\left[{\bf I}-{\bf M}(\epsilon,s)\right]_{ab}^{-1}\widetilde{Q}(|bL-\epsilon|,s).$
@@
${\mathfrak{Q}}_{k_{0}}(x_{0},s)$ $:=\int_{0}^{L}\sum_{k=\pm 1}\widetilde{\rho}_{k|k_{0}}(x,s|x_{0})dx,$ (2.4gsajafgouafa)
$\widetilde{Q}(x_{0},s)$ $:=\int_{0}^{L}\widetilde{p}(x,s|x_{0})dx=\frac{1-\widetilde{J}(L,s|x_{0})+\widetilde{J}(0,s|x_{0})}{s}.$ (2.4gsajafgouafb)
Equation (LABEL:ren2gateQ) is the 2-gate analogue of equation (2.4gsae).
@@
$G(x,s|x_{0})$ $=\frac{\sinh\sqrt{s/D}x\,\sinh\sqrt{s/D}(L-x_{0})}{\sqrt{sD}\sinh\sqrt{s/D}L}\Theta(x_{0}-x)$ (2.4gsajafgouafag)
$\quad+\frac{\sinh\sqrt{s/D}x_{0}\,\sinh\sqrt{s/D}(L-x)}{\sqrt{sD}\sinh\sqrt{s/D}L}\Theta(x-x_{0}),$ (2.4gsajafgouafah)
@@
$\widetilde{J}(0,s|x_{0})$ $=-\frac{\sinh\sqrt{s/D}(L-x_{0})}{\sinh\sqrt{s/D}L},\quad\widetilde{J}(L,s|x_{0})=\frac{\sinh\sqrt{s/D}x_{0}}{\sinh\sqrt{s/D}L}.$ (2.4gsajafgouafai)
Substituting into the matrix (2.4gsajafgouac) and expanding to $O(\epsilon)$ gives
${\bf M}(\epsilon,s)={\bf I}-\epsilon{\bm{\Lambda}}(s)+O(\epsilon^{2}),$ (2.4gsajafgouafaj)
@@
${\bf m}(s)=\frac{\sqrt{{s}/{D}}}{2\sinh\sqrt{s/D}L}\left(\begin{array}[]{cc}\mbox{cosh}\sqrt{s/D}L&-1\\ -1&\mbox{cosh}\sqrt{s/D}L\end{array}\right).$ (2.4gsajafgouafam)
@@
$\widetilde{Q}(L-\epsilon,s)=\widetilde{Q}(\epsilon,s)=\epsilon\widetilde{q}(s)+O(\epsilon^{2}),\quad\widetilde{q}(s)\equiv\frac{\cosh\sqrt{s/D}L-1}{\sqrt{sD}\sinh\sqrt{s/D}L},$ (2.4gsajafgouafan)